# Demystifying amortized causal discovery with transformers

### v2

@@ E.1 Proof of Proposition 1
Under the hypothesis that equations (5), (6) hold, i.e. when the data generating process satisfy both a forward and a backward model, by Theorem 2 we have that:
$\xi^{\prime\prime\prime}(\tilde{x})=\xi^{\prime\prime}(\tilde{x})G(\tilde{x},\tilde{y})+H(\tilde{x},\tilde{y}),$ (15)
@@
Define $z\coloneqq\xi^{\prime\prime\prime}$, such that the above equation can be written as $z^{\prime}(\tilde{x})=z(\tilde{x})G(\tilde{x},\tilde{y})+H(\tilde{x},\tilde{y})$. given that such function $z$ exists, it is given by:
@@
Let $\tilde{y}$ such that $\nu^{\prime\prime}_{Y}(\tilde{y}-h_{Y}(\tilde{x}))h^{\prime}_{Y}(\tilde{x})\neq 0$ holds for all but countable values of $\tilde{x}$. Then, $z$ is determined by $z(\tilde{x}_{0})$, as we can extend equation (16) to all the remaining points. The set of all functions $\xi$ satisfying the differential equation (15) is a 3-dimensional affine space, as fixing $\xi(\tilde{x}_{0}),\xi^{\prime\prime}(\tilde{x}_{0}),\xi^{\prime\prime}(\tilde{x}_{0})$ for some point $\tilde{x}_{0}$ completely determines the solution $\xi$. Moreover, given $\nu_{Y},h_{X},h_{Y}$ fixed, $\xi^{\prime\prime}$ is specified by (10) of theorem 2, which implies:
@@
which confines $\xi$ solutions of (15) to a 2-dimensional affine space. ∎

### v3

@@ E.1 Proof of Proposition 1
Under the hypothesis that equations equation 5, equation 6 hold, i.e. when the data generating process satisfy both a forward and a backward model, by Theorem 2 we have that:
[…]
@@
$z(\tilde{x})=z(\tilde{x}_{0})e^{\int_{\tilde{x}_{0}}^{\tilde{x}}G(t,y)dt}+\int_{\tilde{x}_{0}}^{\tilde{x}}e^{\int_{\hat{t}}^{\tilde{x}}G(t,y)dt}H(\hat{t},y)d\hat{t}$. (16)
Let $\tilde{y}$ such that $\nu^{\prime\prime}_{Y}(\tilde{y}-h_{Y}(\tilde{x}))h^{\prime}_{Y}(\tilde{x})\neq 0$ holds for all but countable values of $\tilde{x}$. Then, $z$ is determined by $z(\tilde{x}_{0})$, as we can extend equation equation 16 to all the remaining points. The set of all functions $\xi$ satisfying the differential equation equation 15 is a 3-dimensional affine space, as fixing $\xi(\tilde{x}_{0}),\xi^{\prime\prime}(\tilde{x}_{0}),\xi^{\prime\prime}(\tilde{x}_{0})$ for some point $\tilde{x}_{0}$ completely determines the solution $\xi$. Moreover, given $\nu_{Y},h_{X},h_{Y}$ fixed, $\xi^{\prime\prime}$ is specified by equation 10 of theorem 2, which implies:
@@
which confines $\xi$ solutions of equation 15 to a 2-dimensional affine space. ∎